# Spectral selectors on lens spaces and applications to the geometry of the group of contactomorphisms

### v1

@@ 3. Spectral selectors
with the infimum taken over all contact Hamiltonian functions $H_{t}$ whose flow represents $\widetilde{\phi}$, is continuous with respect to the $\mathcal{C}^{1}$-topology. Indeed, this can be seen as follows. Since $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ is a topological group, it is enough to show that $\nu_{\alpha}$ is $\mathcal{C}^{1}$-continuous at the identity, i.e. that for every $\epsilon>0$ there is a $\mathcal{C}^{1}$-neighborhood $\widetilde{\mathcal{U}}$ of $\widetilde{\mathrm{id}}$ in $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ such that for every $\widetilde{\phi}\in\widetilde{\mathcal{U}}$ we have $\nu_{\alpha}(\widetilde{\phi})<\epsilon$. As in the proof of Section 3 we consider the product $L_{k}^{2n-1}\times L_{k}^{2n-1}\times\mathbb{R}$ endowed with the contact structure given by the kernel of the contact form $\pi_{2}^{\ast}\alpha_{0}-e^{\theta}\pi_{1}^{\ast}\alpha_{0}$. Applying the Weinstein theorem we can find a neighborhood $\mathcal{U}(\Delta\times\{0\})$ of the Legendrian $\Delta\times\{0\}\cong L_{k}^{2n-1}$ of $L_{k}^{2n-1}\times L_{k}^{2n-1}\times\mathbb{R}$, a neighborhood $\mathcal{U}(j^{1}0)$ of the zero section of $J^{1}L_{k}^{2n-1}$ of the form $\mathcal{U}(j^{1}0)=\underline{\mathcal{U}}(j^{1}0)\times(-\epsilon^{\prime},\epsilon^{\prime})$ with $\epsilon^{\prime}<\epsilon$ and a diffeomorphism $\Psi$ from $\mathcal{U}(\Delta\times\{0\})$ to $\mathcal{U}(j^{1}0)$ with $\Psi(\Delta\times\{0\})=j^{1}0$ and $\Psi^{\ast}(dz-\lambda_{\operatorname{can}})=\pi_{2}^{\ast}\alpha_{0}-e^{\theta}\pi_{1}^{\ast}\alpha_{0}$. Since the map $j^{1}:\mathcal{C}^{\infty}(L_{k}^{2n-1})\rightarrow\mathcal{L}eg(j^{1}0)$ that associates to a function its 1-jet is a local homeomorphism with respect to the $\mathcal{C}^{2}$-topology on $\mathcal{C}^{\infty}(L_{k}^{2n-1})$ and the $\mathcal{C}^{1}$-topology on $\mathcal{L}eg(j^{1}0)$, we can find a convex $\mathcal{C}^{2}$-neighborhood $\mathcal{U}(0)$ of the zero function in $\mathcal{C}^{\infty}(L_{k}^{2n-1})$ such that $j^{1}f\in\mathcal{U}(j^{1}0)$ for any $f\in\mathcal{U}(0)$, and the map $\mathcal{U}(0)\mapsto\mathcal{L}eg(\Delta\times\{0\})$ that sends $f$ to $\Psi^{-1}(j^{1}f)$ is a local homeomorphism. Since $\Delta\times\{0\}=\operatorname{gr}_{\alpha_{0}}(\mathrm{id})$ and the map $\operatorname{gr}_{\alpha_{0}}:\operatorname{Cont}_{0}(L_{k}^{2n-1},\xi_{0})\rightarrow\mathcal{L}eg\big{(}\operatorname{gr}_{\alpha_{0}}(\mathrm{id})\big{)}$ that associates to a contactomorphism its graph is a local homeomorphism with respect to the $\mathcal{C}^{1}$-topologies, we obtain a map $\mathcal{U}(0)\rightarrow\operatorname{Cont}_{0}(L_{k}^{2n-1},\xi_{0})$ that associates to a function $f$ a contactomorphism $\phi$ with $j^{1}f=\Psi(\operatorname{gr}_{\alpha_{0}}(\phi))$, which is a homeomorphism on its image $\mathcal{U}$. Since $\mathcal{U}(0)$ is convex, $\mathcal{U}$ is simply connected; let thus $\widetilde{\mathcal{U}}$ be the open neighborhood of $\widetilde{\mathrm{id}}$ in $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ that projects homeomorphically to $\mathcal{U}$. Consider now any $\widetilde{\phi}$ in $\widetilde{\mathcal{U}}$. Let $f:L_{k}^{2n-1}\rightarrow(-\epsilon,\epsilon)$ be the function in $\mathcal{U}(0)$ such that $j^{1}f=\Psi\big{(}\operatorname{gr}_{\alpha_{0}}\big{(}\Pi(\widetilde{\phi})\big{)}\big{)}$. Since $\mathcal{U}(0)$ is convex, $f_{t}:=tf$ is in $\mathcal{U}(0)$ for every $t\in[0,1]$, and so for every $t\in[0,1]$ there is $\phi_{t}\in\mathcal{U}$ with $j^{1}f_{t}=\Psi\big{(}\operatorname{gr}_{\alpha_{0}}(\phi_{t})\big{)}$. Consider the two Legendrian isotopies $j^{1}f_{t}$ and $\operatorname{gr}_{\alpha_{0}}(\phi_{t})$, with parametrizations given respectively by
$i_{1}:[0,1]\times j^{1}0\rightarrow J^{1}L_{k}^{2n-1}\,,\;i_{1}\big{(}t,(x,0,0)\big{)}=j^{1}f_{t}(x)$
@@
$i_{2}:[0,1]\times\big{(}\Delta\times\{0\}\big{)}\rightarrow L_{k}^{2n-1}\times L_{k}^{2n-1}\times\mathbb{R}\,,\;i_{2}\big{(}t,(x,x,0)\big{)}=\big{(}x,\phi_{t}(x),g_{t}(x)\big{)}\,,$
@@
and so $\lvert H_{t}\lvert<\epsilon$. Moreover [$[\{\phi_{t}\}_{t\in[0,1]}]=\widetilde{\phi}$, because $\widetilde{\phi}\cdot[\{\phi_{t}\}_{t\in[0,1]}]^{-1}\in\pi_{1}\big{(}\operatorname{Cont}_{0}(L_{k}^{2n-1},\xi_{0})\big{)}$ can be represented by a loop in $\mathcal{U}$, which is simply connected. We thus conclude that $\nu_{\alpha}(\widetilde{\phi})<\epsilon$, as we wanted.
Using (11) and the fact that $\nu_{\alpha}$ is $\mathcal{C}^{1}$-continuous we now deduce that each $c_{j}$ is $\mathcal{C}^{1}$-continuous. Let $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$. By $\mathcal{C}^{1}$-continuity of $\nu_{\alpha}$, for any $\epsilon>0$ there is a $\mathcal{C}^{1}$-neighborhood $\widetilde{\mathcal{U}}$ of $\widetilde{\mathrm{id}}$ in $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ such that $\left.\nu_{\alpha}\right\lvert_{\widetilde{\mathcal{U}}}<\epsilon$. Then $\widetilde{\mathcal{V}}:=\widetilde{\mathcal{U}}\cdot\widetilde{\phi}$ is a $\mathcal{C}^{1}$-neighborhood of $\widetilde{\phi}$ such that for every $\widetilde{\psi}\in\widetilde{\mathcal{V}}$ we have
@@
$c_{j+l}(\widetilde{\phi}\cdot\widetilde{\psi})\leq c_{j}(\widetilde{\phi})+\left\lceil c_{l}(\widetilde{\psi})\right\rceil_{\frac{2\pi}{k}}\,.$
Let $T=c_{j}(\widetilde{\phi})$ and $N=\lceil c_{l}(\widetilde{\psi})\rceil_{\frac{2\pi}{k}}$. Since $\Pi:\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\rightarrow\operatorname{Cont}_{0}(L_{k}^{2n-1},\xi_{0})$ is a local homeomorphism, by the continuity property of the spectral selectors and Section 3 we can assume that $\Pi(\widetilde{\phi})$ is non-degenerate with respect to $\alpha_{0}$. By the relation with translated points property, the spectral selectors of $\widetilde{\phi}$ are thus all distinct, and so $\mu\big{(}\widetilde{r_{-T}}\cdot\widetilde{\phi}\big{)}=-j$. Using the fact that $\widetilde{r_{-N}}$ commutes with $\widetilde{\phi}$ (since $\{r_{t}\}$ is $\frac{2\pi}{k}$-periodic) and the triangle inequality for the nonlinear Maslov index (Section 2 (iii)) we thus have
@@
$c_{j+l}\big{(}\widetilde{\phi}\cdot\widetilde{\psi}\big{)}\leq T+N=c_{j}(\widetilde{\phi})+\left\lceil c_{l}(\widetilde{\psi})\right\rceil_{\frac{2\pi}{k}}\,.$
@@
$\left\lceil c_{j}(\widetilde{\psi}\cdot\widetilde{\phi}\cdot\widetilde{\psi}^{-1})\right\rceil_{\frac{2\pi}{k}}=\left\lceil c_{j}(\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}}\,.$ (12)
@@
is continuous. Moreover, $c_{j}(\widetilde{\psi}_{s}\cdot\widetilde{\phi}\cdot\widetilde{\psi}_{s}^{-1})\in\mathbb{R}\smallsetminus\frac{2\pi}{k}\cdot\mathbb{Z}$ for all $s\in[0,1]$. Indeed, if we had $c_{j}(\widetilde{\psi}_{\underline{s}}\cdot\widetilde{\phi}\cdot\widetilde{\psi}_{\underline{s}}^{-1})\in\frac{2\pi}{k}\cdot\mathbb{Z}$ for some $\underline{s}$ then, by the spectrality property, $\Pi(\widetilde{\psi}_{\underline{s}}\cdot\widetilde{\phi}\cdot\widetilde{\psi}_{\underline{s}}^{-1})$ would have discriminant points. But this is absurd, because the discriminant points of $\Pi(\widetilde{\psi}_{\underline{s}}\cdot\widetilde{\phi}\cdot\widetilde{\psi}_{\underline{s}}^{-1})$ are in bijection with the discriminant points of $\Pi(\widetilde{\phi})$. We thus obtain (12) in this case.
The general case can be obtained as follows. Given any $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$, since (by Section 3) $\mathcal{A}(\widetilde{\phi})$ is nowhere dense, there is a sequence $(\epsilon_{n})$ of positive real numbers with $\epsilon_{n}\to 0$ such that, for every $n$, $\Pi(\widetilde{r_{-\epsilon_{n}}}\cdot\widetilde{\phi})$ does not have discriminant points. Pose $\widetilde{\chi_{n}}=\widetilde{r_{-\epsilon_{n}}}\cdot\widetilde{\phi}$. By the first part of the proof we have
$\left\lceil c_{j}(\widetilde{\psi}\cdot\widetilde{\chi_{n}}\cdot\widetilde{\psi}^{-1})\right\rceil_{\frac{2\pi}{k}}=\Big{\lceil}c_{j}(\widetilde{\chi}_{n})\Big{\rceil}_{\frac{2\pi}{k}}$ (13)
for all $n$. Since $(\widetilde{\chi}_{n})$ converges to $\widetilde{\phi}$ in the $\mathcal{C}^{1}$-topology and $\widetilde{\chi}_{n}\leq\widetilde{\phi}$ for all $n$, by the continuity and monotonicity properties of the spectral selectors for $n$ big enough we have $\left\lceil c_{j}(\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}}=\left\lceil c_{j}(\widetilde{\chi}_{n})\right\rceil_{\frac{2\pi}{k}}$ and $\left\lceil c_{j}(\widetilde{\psi}\cdot\widetilde{\phi}\cdot\widetilde{\psi}^{-1})\right\rceil_{\frac{2\pi}{k}}=\left\lceil c_{j}(\widetilde{\psi}\cdot\widetilde{\chi}_{n}\cdot\widetilde{\psi}^{-1})\right\rceil_{\frac{2\pi}{k}}$. Equation (13) thus gives the desired result (12).
@@
$\left\lfloor c_{j}\big{(}\widetilde{\phi}\big{)}\right\rfloor_{\frac{2\pi}{k}}=\max\left\{\,N\in\frac{2\pi}{k}\cdot\mathbb{Z}\;\ \Big{\lvert}\;\mu(\widetilde{r_{-N}}\cdot\widetilde{\phi})>-j\,\right\}\,.$ (14)
Indeed, by spectrality we have $c_{j}(\widetilde{\phi})\notin\frac{2\pi}{k}\cdot\mathbb{Z}$ and thus $\underline{N}:=\big{\lfloor}c_{j}\big{(}\widetilde{\phi}\big{)}\big{\rfloor}_{\frac{2\pi}{k}}<c_{j}(\widetilde{\phi})$. This implies that $\mu(\widetilde{r_{-\underline{N}}}\cdot\widetilde{\phi})>-j$, and so the inequality $\leq$ in (14). On the other hand, the opposite inequality follows (without any assumption on $\widetilde{\phi}$) from (10) and the fact that $\big{\lfloor}c_{j}\big{(}\widetilde{\phi}\big{)}\big{\rfloor}_{\frac{2\pi}{k}}+\frac{2\pi}{k}\geq\big{\lceil}c_{j}\big{(}\widetilde{\phi}\big{)}\big{\rceil}_{\frac{2\pi}{k}}$.
@@
$\left\lceil c_{j}(\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}}=-\left\lfloor c_{-j-(2n-1)}(\widetilde{\phi}^{-1})\right\rfloor_{\frac{2\pi}{k}}$ (15)
for any $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$. Assume first that $\Pi\big{(}\mathcal{\widetilde{\phi}}\big{)}$ does not have discriminant points. Then Section 2 (ix) implies that
@@
for every $N$ that is a multiple of $\frac{2\pi}{k}$. The Poincaré duality (15) then follows from (10), (14), (16) and the fact that $\widetilde{\phi}^{-1}\cdot\widetilde{r_{-N}}=\widetilde{r_{-N}}\cdot\widetilde{\phi}^{-1}$ for every $N$ that is a multiple of $\frac{2\pi}{k}$.
For a general $\widetilde{\phi}$, as in the proof of conjugation invariance we can find a sequence $(\widetilde{\chi}_{n})$ that converges to $\widetilde{\phi}$ in the $\mathcal{C}^{1}$-topology and such that, for all $n$, $\widetilde{\chi}_{n}\leq\widetilde{\phi}$ and $\Pi(\widetilde{\chi}_{n})$ does not have discriminant points. By the first part of the proof we have $\left\lceil c_{j}(\widetilde{\chi}_{n})\right\rceil_{\frac{2\pi}{k}}=-\left\lfloor c_{-j-(2n-1)}(\widetilde{\chi}_{n}^{-1})\right\rfloor_{\frac{2\pi}{k}}$. By monotonicity and continuity of the spectral selectors we thus obtain (15) also in this case.

### v2

@@ 3. Spectral selectors
with the infimum taken over all contact Hamiltonian functions $H_{t}$ whose flow represents $\widetilde{\phi}$, is continuous with respect to the $\mathcal{C}^{1}$-topology. This follows from the main result of [5], or can be seen directly as follows. Since $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ is a topological group, it is enough to show that $\nu_{\alpha}$ is $\mathcal{C}^{1}$-continuous at the identity, i.e. that for every $\epsilon>0$ there is a $\mathcal{C}^{1}$-neighborhood $\widetilde{\mathcal{U}}$ of $\widetilde{\mathrm{id}}$ in $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ such that for every $\widetilde{\phi}\in\widetilde{\mathcal{U}}$ we have $\nu_{\alpha}(\widetilde{\phi})<\epsilon$. As in the proof of Section 3 we consider the product $L_{k}^{2n-1}(\underline{w})\times L_{k}^{2n-1}(\underline{w})\times\mathbb{R}$ endowed with the contact structure given by the kernel of the contact form $\pi_{2}^{\ast}\alpha_{0}-e^{\theta}\pi_{1}^{\ast}\alpha_{0}$. Applying the Weinstein theorem we can find a neighborhood $\mathcal{U}(\operatorname{gr}_{\alpha_{0}}(\mathrm{id}))$ of $\operatorname{gr}_{\alpha_{0}}(\mathrm{id})$, a neighborhood $\mathcal{U}(j^{1}0)$ of the zero section of $J^{1}L_{k}^{2n-1}(\underline{w})$ of the form $\mathcal{U}(j^{1}0)=\underline{\mathcal{U}}(j^{1}0)\times(-\epsilon^{\prime},\epsilon^{\prime})$ with $\epsilon^{\prime}<\epsilon$ and a diffeomorphism $\Psi$ from $\mathcal{U}(\operatorname{gr}_{\alpha_{0}}(\mathrm{id}))$ to $\mathcal{U}(j^{1}0)$ with $\Psi(\operatorname{gr}_{\alpha_{0}}(\mathrm{id}))=j^{1}0$ and $\Psi^{\ast}(dz-\lambda_{\operatorname{can}})=\pi_{2}^{\ast}\alpha_{0}-e^{\theta}\pi_{1}^{\ast}\alpha_{0}$. Since the map $j^{1}:\mathcal{C}^{\infty}\big{(}L_{k}^{2n-1}(\underline{w})\big{)})\rightarrow\mathcal{L}eg(j^{1}0)$ that associates to a function its 1-jet is a local homeomorphism with respect to the $\mathcal{C}^{2}$-topology on $\mathcal{C}^{\infty}\big{(}L_{k}^{2n-1}(\underline{w})\big{)}$ and the $\mathcal{C}^{1}$-topology on $\mathcal{L}eg(j^{1}0)$, we can find a convex $\mathcal{C}^{2}$-neighborhood $\mathcal{U}(0)$ of the zero function in $\mathcal{C}^{\infty}\big{(}L_{k}^{2n-1}(\underline{w})\big{)}$ such that $j^{1}f\in\mathcal{U}(j^{1}0)$ for any $f\in\mathcal{U}(0)$, and the map $\mathcal{U}(0)\mapsto\mathcal{L}eg(\operatorname{gr}_{\alpha_{0}}(\mathrm{id}))$ that sends $f$ to $\Psi^{-1}(j^{1}f)$ is a local homeomorphism. Since the map $\operatorname{gr}_{\alpha_{0}}:\operatorname{Cont}_{0}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow\mathcal{L}eg\big{(}\operatorname{gr}_{\alpha_{0}}(\mathrm{id})\big{)}$ that associates to a contactomorphism its graph is a local homeomorphism with respect to the $\mathcal{C}^{1}$-topologies, we obtain a map $\mathcal{U}(0)\rightarrow\operatorname{Cont}_{0}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ that associates to a function $f$ a contactomorphism $\phi$ with $j^{1}f=\Psi(\operatorname{gr}_{\alpha_{0}}(\phi))$, which is a homeomorphism on its image $\mathcal{U}$. Since $\mathcal{U}(0)$ is convex, $\mathcal{U}$ is simply connected; let thus $\widetilde{\mathcal{U}}$ be the open neighborhood of $\widetilde{\mathrm{id}}$ in $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ that projects homeomorphically to $\mathcal{U}$. Consider now any $\widetilde{\phi}$ in $\widetilde{\mathcal{U}}$. Let $f:L_{k}^{2n-1}(\underline{w})\rightarrow(-\epsilon,\epsilon)$ be the function in $\mathcal{U}(0)$ such that $j^{1}f=\Psi\big{(}\operatorname{gr}_{\alpha_{0}}\big{(}\Pi(\widetilde{\phi})\big{)}\big{)}$. Since $\mathcal{U}(0)$ is convex, $f_{t}:=tf$ is in $\mathcal{U}(0)$ for every $t\in[0,1]$, and so for every $t\in[0,1]$ there is $\phi_{t}\in\mathcal{U}$ with $j^{1}f_{t}=\Psi\big{(}\operatorname{gr}_{\alpha_{0}}(\phi_{t})\big{)}$. Consider the two Legendrian isotopies $j^{1}f_{t}$ and $\operatorname{gr}_{\alpha_{0}}(\phi_{t})$, with parametrisations given respectively by
$i_{1}:[0,1]\times j^{1}0\rightarrow J^{1}L_{k}^{2n-1}(\underline{w})\,,\;i_{1}\big{(}t,(x,0,0)\big{)}=j^{1}f_{t}(x)$
@@
$i_{2}:[0,1]\times\operatorname{gr}_{\alpha_{0}}(\mathrm{id})\rightarrow L_{k}^{2n-1}(\underline{w})\times L_{k}^{2n-1}(\underline{w})\times\mathbb{R}\,,\;i_{2}\big{(}t,(x,x,0)\big{)}=\big{(}x,\phi_{t}(x),g_{t}(x)\big{)}\,,$
@@
and so $\lvert H_{t}\lvert<\epsilon$. Moreover [$[\{\phi_{t}\}_{t\in[0,1]}]=\widetilde{\phi}$, because $\widetilde{\phi}\cdot[\{\phi_{t}\}_{t\in[0,1]}]^{-1}$ can be represented by a loop in $\mathcal{U}$, which is simply connected. We thus conclude that $\nu_{\alpha}(\widetilde{\phi})<\epsilon$, as we wanted.
Using (11) and the fact that $\nu_{\alpha}$ is $\mathcal{C}^{1}$-continuous we now deduce that each $c_{j}$ is $\mathcal{C}^{1}$-continuous. Let $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$. By $\mathcal{C}^{1}$-continuity of $\nu_{\alpha}$, for any $\epsilon>0$ there is a $\mathcal{C}^{1}$-neighborhood $\widetilde{\mathcal{U}}$ of $\widetilde{\mathrm{id}}$ in $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ such that $\left.\nu_{\alpha}\right\lvert_{\widetilde{\mathcal{U}}}<\epsilon$. Then $\widetilde{\mathcal{V}}:=\widetilde{\mathcal{U}}\cdot\widetilde{\phi}$ is a $\mathcal{C}^{1}$-neighborhood of $\widetilde{\phi}$ such that for every $\widetilde{\psi}\in\widetilde{\mathcal{V}}$ we have
@@
$c_{j+l}(\widetilde{\phi}\cdot\widetilde{\psi})\leq c_{j}(\widetilde{\phi})+\left\lceil c_{l}(\widetilde{\psi})\right\rceil_{T_{\underline{w}}}\,.$
Let $T=c_{j}(\widetilde{\phi})$ and $N=\lceil c_{l}(\widetilde{\psi})\rceil_{T_{\underline{w}}}$. Since $\Pi:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow\operatorname{Cont}_{0}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ is a local homeomorphism, by the continuity property of the spectral selectors and Section 3 we can assume that $\Pi(\widetilde{\phi})$ is non-degenerate with respect to $\alpha_{0}$. By the relation with translated points property, the spectral selectors of $\widetilde{\phi}$ are thus all distinct, and so $\mu\big{(}\widetilde{r_{-T}}\cdot\widetilde{\phi}\big{)}=-j$. Using the fact that $\widetilde{r_{-N}}$ commutes with $\widetilde{\phi}$ (since $\{r_{t}\}$ has period $T_{\underline{w}}$) and the triangle inequality for the non-linear Maslov index (Section 2 (iii)) we thus have
@@
$c_{j+l}\big{(}\widetilde{\phi}\cdot\widetilde{\psi}\big{)}\leq T+N=c_{j}(\widetilde{\phi})+\left\lceil c_{l}(\widetilde{\psi})\right\rceil_{T_{\underline{w}}}\,.$
@@
$\left\lceil c_{j}(\widetilde{\psi}\cdot\widetilde{\phi}\cdot\widetilde{\psi}^{-1})\right\rceil_{T_{\underline{w}}}=\left\lceil c_{j}(\widetilde{\phi})\right\rceil_{T_{\underline{w}}}\,.$ (12)
@@
is continuous. Moreover, $c_{j}(\widetilde{\psi}_{s}\cdot\widetilde{\phi}\cdot\widetilde{\psi}_{s}^{-1})\in\mathbb{R}\smallsetminus T_{\underline{w}}\cdot\mathbb{Z}$ for all $s\in[0,1]$. Indeed, if we had $c_{j}(\widetilde{\psi}_{\underline{s}}\cdot\widetilde{\phi}\cdot\widetilde{\psi}_{\underline{s}}^{-1})\in T_{\underline{w}}\cdot\mathbb{Z}$ for some $\underline{s}$ then, by the spectrality property, $\Pi(\widetilde{\psi}_{\underline{s}}\cdot\widetilde{\phi}\cdot\widetilde{\psi}_{\underline{s}}^{-1})$ would have discriminant points. But this is absurd, because the discriminant points of $\Pi(\widetilde{\psi}_{\underline{s}}\cdot\widetilde{\phi}\cdot\widetilde{\psi}_{\underline{s}}^{-1})$ are in bijection with the discriminant points of $\Pi(\widetilde{\phi})$. We thus obtain (12) in this case.
The general case can be obtained as follows. Given any $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$, since (by Section 3) $\mathcal{A}(\widetilde{\phi})$ is nowhere dense, there is a sequence $(\epsilon_{l})$ of positive real numbers with $\epsilon_{l}\to 0$ such that, for every $l$, $\Pi(\widetilde{r_{-\epsilon_{l}}}\cdot\widetilde{\phi})$ does not have discriminant points. Pose $\widetilde{\chi_{l}}=\widetilde{r_{-\epsilon_{l}}}\cdot\widetilde{\phi}$. By the first part of the proof we have
$\left\lceil c_{j}(\widetilde{\psi}\cdot\widetilde{\chi_{l}}\cdot\widetilde{\psi}^{-1})\right\rceil_{T_{\underline{w}}}=\Big{\lceil}c_{j}(\widetilde{\chi}_{l})\Big{\rceil}_{T_{\underline{w}}}$ (13)
for all $l$. Since $(\widetilde{\chi}_{l})$ converges to $\widetilde{\phi}$ in the $\mathcal{C}^{1}$-topology and $\widetilde{\chi}_{l}\leq\widetilde{\phi}$ for all $l$, by the continuity and monotonicity properties of the spectral selectors for $l$ big enough we have $\left\lceil c_{j}(\widetilde{\phi})\right\rceil_{T_{\underline{w}}}=\left\lceil c_{j}(\widetilde{\chi}_{l})\right\rceil_{T_{\underline{w}}}$ and $\left\lceil c_{j}(\widetilde{\psi}\cdot\widetilde{\phi}\cdot\widetilde{\psi}^{-1})\right\rceil_{T_{\underline{w}}}=\left\lceil c_{j}(\widetilde{\psi}\cdot\widetilde{\chi}_{l}\cdot\widetilde{\psi}^{-1})\right\rceil_{T_{\underline{w}}}$. Equation (13) thus gives the desired result (12).
@@
$\left\lfloor c_{j}\big{(}\widetilde{\phi}\big{)}\right\rfloor_{T_{\underline{w}}}=\max\left\{\,N\in T_{\underline{w}}\cdot\mathbb{Z}\;\ \Big{\lvert}\;\mu(\widetilde{r_{-N}}\cdot\widetilde{\phi})>-j\,\right\}\,.$ (14)
Indeed, by spectrality we have $c_{j}(\widetilde{\phi})\notin T_{\underline{w}}\cdot\mathbb{Z}$ and thus $\underline{N}:=\big{\lfloor}c_{j}\big{(}\widetilde{\phi}\big{)}\big{\rfloor}_{T_{\underline{w}}}<c_{j}(\widetilde{\phi})$. This implies that $\mu(\widetilde{r_{-\underline{N}}}\cdot\widetilde{\phi})>-j$, and so the inequality $\leq$ in (14). On the other hand, the opposite inequality follows (without any assumption on $\widetilde{\phi}$) from (10) and the fact that $\big{\lfloor}c_{j}\big{(}\widetilde{\phi}\big{)}\big{\rfloor}_{T_{\underline{w}}}+T_{\underline{w}}\geq\big{\lceil}c_{j}\big{(}\widetilde{\phi}\big{)}\big{\rceil}_{T_{\underline{w}}}$.
@@
$\left\lceil c_{j}(\widetilde{\phi})\right\rceil_{T_{\underline{w}}}=-\left\lfloor c_{-j-(2n-1)}(\widetilde{\phi}^{-1})\right\rfloor_{T_{\underline{w}}}$ (15)
for any $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$. Assume first that $\Pi\big{(}\mathcal{\widetilde{\phi}}\big{)}$ does not have discriminant points. Then Section 2 (ix) implies that
@@
for every $N$ that is a multiple of $T_{\underline{w}}$. The Poincaré duality (15) then follows from (10), (14), (16) and the fact that $\widetilde{\phi}^{-1}\cdot\widetilde{r_{-N}}=\widetilde{r_{-N}}\cdot\widetilde{\phi}^{-1}$ for every $N$ that is a multiple of $T_{\underline{w}}$.
For a general $\widetilde{\phi}$, as in the proof of conjugation invariance we can find a sequence $(\widetilde{\chi}_{l})$ that converges to $\widetilde{\phi}$ in the $\mathcal{C}^{1}$-topology and such that, for all $l$, $\widetilde{\chi}_{l}\leq\widetilde{\phi}$ and $\Pi(\widetilde{\chi}_{l})$ does not have discriminant points. By the first part of the proof we have $\left\lceil c_{j}(\widetilde{\chi}_{l})\right\rceil_{T_{\underline{w}}}=-\left\lfloor c_{-j-(2n-1)}(\widetilde{\chi}_{l}^{-1})\right\rfloor_{T_{\underline{w}}}$. By monotonicity and continuity of the spectral selectors we thus obtain (15) also in this case.